# Bosch Research North America & Bosch Center for AI https://zixunh.github.io/3d-geer

### v1

@@ Abstract.
3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) marks a significant milestone in balancing the quality and efficiency of differentiable rendering. However, its high efficiency stems from an approximation of projecting 3D Gaussians onto the image plane as 2D Gaussians, which inherently limits rendering quality—particularly under large Field-of-View (FoV) camera inputs. While several recent works have extended 3DGS to mitigate these approximation errors, none have successfully achieved both exactness and high efficiency simultaneously. In this work, we introduce 3DGEER , an Exact and Efficient Volumetric Gaussian Rendering method. Starting from first principles, we derive a closed-form expression for the density integral along a ray traversing a 3D Gaussian distribution. This formulation enables precise forward rendering with arbitrary camera models and supports gradient-based optimization of 3D Gaussian parameters. To ensure both exactness and real-time performance, we propose an efficient method for computing a tight Particle Bounding Frustum (PBF) for each 3D Gaussian, enabling accurate and efficient ray-Gaussian association. We also introduce a novel Bipolar Equiangular Projection (BEAP) representation to accelerate ray association under generic camera models. BEAP further provides a more uniform ray sampling strategy to apply supervision, which empirically improves reconstruction quality. Experiments on multiple pinhole and fisheye datasets show that our method consistently outperforms prior methods, establishing a new state-of-the-art in real-time neural rendering.
Volumetric Rendering, Differentiable Rendering, Novel View Synthesis, Radiance Fields, Neural Reconstruction
journal: TOGccs: Computing methodologies Renderingccs: Computing methodologies Reconstructionccs: Computing methodologies Machine Learning Approaches22footnotetext: Corresponding Author.
1. Introduction
@@
Initiated by 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023), volumetric particle rendering methods formulated under splatting (Yu et al., 2024a; Yan et al., 2024; Lu et al., 2024; Kheradmand et al., 2024; Charatan et al., 2024; Mallick et al., 2024) have gained significant popularity due to their impressive visual fidelity and high rendering speed. The core efficiency stems from EWA splatting (Zwicker et al., 2001), which uses a local affine approximation to project a 3D Gaussian as a 2D Gaussian on the image plane. While effective in narrow field-of-view (FoV) scenarios, this approximation introduces substantial errors in wide-FoV settings, where nonlinear distortions cause the true projection to deviate significantly from a symmetric 2D Gaussian—especially in fisheye images commonly encountered in robotics and autonomous driving (Rashed et al., 2021; Courbon et al., 2007; Kumar et al., 2023; Zhang et al., 2016; Yogamani et al., 2019, 2024; Sekkat et al., 2022). Although recent works have attempted to enable differentiable rendering under specific fisheye camera models (Liao et al., 2024) or more general camera models (Huang et al., 2024), these approaches remain fundamentally constrained by their reliance on 2D projective-space approximations and have yet to yield an exact closed-form solution.
@@
A parallel line of work tackles volumetric particle rendering directly in a ray-marching formulation (Yu et al., 2024b; Condor et al., 2025; Mai et al., 2024; Moënne-Loccoz et al., 2024). These approaches naturally support large-FoV camera models and are free from projective approximation error. However, achieving high efficiency remains a key challenge—particularly in associating each rendering ray with its most relevant 3D particles. To address this, significant effort has been devoted to accelerating it, such as adopting advanced Bounding Volume Hierarchies (BVH) employed in EVER (Mai et al., 2024) and 3DGRT (Moënne-Loccoz et al., 2024). Nevertheless, due to the inherent algorithmic complexity and difficulty in parallelization, pure ray marching methods have yet to match the efficiency of 3DGS. Alternatively, a hybrid approach has been proposed in 3DGUT (Wu et al., 2025) which adopts ray-tracing rendering with rasterization to solve ray-particle association efficiently. However, the inherited projective approximation error from association stage still compromises rendering exactness. To date, no existing method achieves both the exactness of ray-based rendering and the high efficiency of projection-based splatting.
@@
In this work, we propose 3DGEER (3D Gaussian Exact and Efficient Rendering)—a novel volumetric Gaussian rendering method that achieves exact rendering while maintaining high efficiency comparable to 3DGS (see Fig. 1). 3DGEER is built upon three key technical contributions: (i) Instead of relying on projective approximations or 2D conic fitting, we revisit ray-marching-based volumetric Gaussian rendering from first principles and derive a closed-form solution for integrating density along a 3D ray through a Gaussian. Based on it, we further derive the exact volumetric Gaussian rendering framework. (ii) To perform efficient ray-particle association without compromising rendering exactness, we reformulate the problem as associating Camera Sub-frustums (CSF) with Particle Bounding Frustums (PBF)—analogous to the tile-AABB mapping used in 3DGS. For each 3D Gaussian, we derive a closed-form solution to compute the PBF tightly bounding it, enabling highly efficient association. (iii) In addition, we propose to uniformly sample rays in a novel Bipolar Equiangular Projection (BEAP) space to apply color supervision. This ray sampling strategy not only aligns image-space partitioning and the underlying CSF but also improved conventional pinhole or equidistant projections in reconstruction quality. Extensive experiments on both fisheye and pinhole datasets—including ScanNet++ (Yeshwanth et al., 2023), Zip-NeRF (Barron et al., 2023), and MipNeRF (Barron et al., 2021)—show that 3DGEER consistently outperforms prior and concurrent methods, achieving 0.9–4.8 PSNR gains under challenging wide-FoV conditions, and 0.3–2.0 on pinhole views.
In summary, our contributions are as follows:
  1. (1)
    We present the first complete, first-principle derived, closed-form solution for exact volumetric Gaussian rendering;
  2. (2)
    We propose an exact and efficient ray-particle association method that enables high speed rendering;
  3. (3)
    We introduce an equiangular ray sampling strategy that provides more spatially uniform color supervision.
  4. (4)
    Our method consistently outperforms SOTA approaches across multiple datasets, with notable gains on wide-FoV neural reconstruction benchmarks.
@@
Table 1. Summary of Volumetric Particle Rendering Methods wrt. Exactness. The top section lists splatting-based methods, while the bottom section presents ray-marching-based approaches—some of which still rely on projective approximations (e.g., EWA or Unscented Transform) during ray-particle association.
@@
Meth. Scope Type FoV FPS Meth.
3DGS (2023) Splat. EWA Img. 3D Gauss.
FisheyeGS (2024) Splat. EWA Img. 3D Gauss.
GS++  (2024) Splat. EWA Img. 3D Gauss.
GOF (2024b) Ray EWA Img. 3D Gauss.
EVER (2024) Ray BVH Scene 3D Ellip.
3DGRT (2024) Ray BVH Scene 3D Gauss.
3DGUT (2025) Ray UT Img. 3D Gauss.
@@
2. Related Works
2.1. Splatting-Based Gaussian Rendering
Most 3DGS variants  (Yu et al., 2024a; Yan et al., 2024; Lu et al., 2024; Kheradmand et al., 2024; Zhang et al., 2024; Rota Bulò et al., 2024; Chen et al., 2024b; Lee et al., 2024a; Kerbl et al., 2024; Guédon and Lepetit, 2024; Chen et al., 2024a; Chung et al., 2024; Lee et al., 2024b; Charatan et al., 2024; Mallick et al., 2024; Cheng et al., 2024) adopt the original EWA splatting (Zwicker et al., 2001) strategy, projecting 3D Gaussians onto the image plane and approximating their appearance as 2D conics. To support differentiable rendering under large-FoV camera models with distortion, methods such as FisheyeGS (Liao et al., 2024) and GS++ (Huang et al., 2024) estimate the Jacobian of equidistant or spherical projections to adjust the 2D mean and covariance of the projected Gaussians. However, these splatting-based approaches still rely on a first-order Taylor expansion of the highly non-linear projection function. In wide-FoV settings, higher-order terms is non-negligible, resulting in significant approximation errors and degraded reconstruction quality.
@@ 2.1. Splatting-Based Gaussian Rendering
2.2. Ray-Marching-Based Gaussian Rendering
A few recent works have approached volumetric particle rendering using a ray-marching formulation, which inherently supports arbitrary camera models without introducing projective approximation error. 3DGRT (Moenne-Loccoz et al., 2024), inspired by fuzzy metaball rendering (Keselman and Hebert, 2022), approximates the ray-Gaussian interaction as the location of maximum intensity response. This heuristic enables ray-space sampling without explicit projection, effectively mitigating wide-FoV distortion. However, it incurs significant computational overhead due to the high cost of ray-particle association via scene-level spatial partitioning structures such as BVH, ultimately resulting in low frame rates.
Alternatively, EVER (Mai et al., 2024) replaces the Gaussian representation with simplified ellipsoids of constant density, enabling a closed-form solution for transmittance-weighted ray integration. However, it reduces the expressive capacity of the representation compared to full 3D Gaussians, and continues to suffer from low runtime efficiency due to its reliance on BVH-based structures.
Figure 2. Projective Approximation Error in Ray-Particle Association. When projecting a Gaussian particle from 3D space onto the image plane, both EWA and Unscented Transform (UT)-based association methods approximate the 2D covariance of the projection. They then either tighten the bounding box—risking mismatched culling—or overestimate it using the maximum projected radius of the conic. In contrast, our PBF method avoids this intermediate approximation by directly computing the exact bounding structure from the true 3D covariance.
@@ 2.2. Ray-Marching-Based Gaussian Rendering
3DGUT (Wu et al., 2025), a follow-up to 3DGRT, improves runtime efficiency by adopting splatting-based particle association, achieving frame rates comparable to 3DGS. However, its acceleration relies on the Unscented Transform (UT), a sampling-based approximation used to estimate 2D conics in projective image space. While the rendering stage is ray-based, the ray-particle association still introduces approximation errors akin to those in EWA splatting. A similar limitation exists in GOF (Yu et al., 2024b), which directly applies the EWA trick to fit 2D bounding conics and projective centers for association. Notably, under extreme FoV conditions, imprecise AABB estimation and the resulting inaccurate tile-to-AABB associations can lead to visible rendering artifacts (see Figure 2).
@@
In contrast to these methods, our approach solves the ray-particle association problem directly at the frustum level in 3D space, preserving the exactness of ray-based rendering across arbitrary camera models while maintaining rendering speeds comparable to 3DGS. Table 1 summarizes representative methods that aim to address projective approximation errors, highlighting key differences across several important dimensions.
3. 3DGEER
Our method, 3DGEER, is built upon three key components: an exact ray-based volumetric Gaussian rendering formulation, an efficient frustum-particle association strategy, and an effective image representation. In Sec.3.1, we introduce the necessary preliminaries. Sec.3.2 presents the closed-form solution for Gaussian density integral along a 3D ray, followed by our differentiable volumetric rendering method that supports arbitrary FoV camera models. In Sec.3.3, we propose a novel and efficient approach for ray-particle association based on fast, exact computation of the Particle Bounding Frustum (PBF). Finally, in Sec.3.4, we detail the Bipolar Equiangular Projection (BEAP) image representation, which seamlessly supports generic camera models and accelerates frustum-particle association while improving reconstruction quality in practice.
3.1. Preliminaries
We follow the classic parameterization of 3D Gaussian primitives as introduced in prior work (Zwicker et al., 2001). Each 3D Gaussian distribution is defined by a mean position $\boldsymbol{\mu}\in\mathbb{R}^{3}$ and a 3D covariance matrix $\Sigma$ in the world coordinate system:
(1) $\mathcal{G}_{\Sigma,\boldsymbol{\mu}}(\mathbf{x})=\frac{1}{\rho\left|\Sigma\right|^{1/2}}\exp\left(-\frac{1}{2}(\mathbf{x}-\boldsymbol{\mu})^{\top}\Sigma^{-1}(\mathbf{x}-\boldsymbol{\mu})\right),$
[…]
(3) $C(\mathbf{o},\mathbf{d})=\sum_{i}c_{i}(\mathbf{o},\boldsymbol{\mu}_{i})T_{i}\prod_{j=1}^{i-1}(1-T_{j}),$
where $c_{i}(\mathbf{o},\boldsymbol{\mu}_{i})$ derives the view-dependent color of the $i$-th Gaussian as seen from the optical center $\mathbf{o}$, parameterized using spherical harmonics (SH) (Kerbl et al., 2023). Through minimizing the reconstruction loss between the rendered color and the observed color along each ray, the parameters $\mathbf{s},\mathbf{q},\boldsymbol{\mu}$, the coefficient $\sigma$ as well as the SH coefficients, are jointly optimized via backpropagation.
3.2. Exact Volumetric Gaussian Rendering
The core challenge in achieving exact volumetric Gaussian rendering in a ray marching formulation lies in obtaining a closed-form solution for integrating Gaussian density and color along a 3D ray. To address this, we first transform the ray—relative to each Gaussian—into a canonical coordinate system where the transmittance integral (see Eq. B.4) remains consistent and the Gaussians become isotropic. We then derive the exact ray color from first principles.
(4) $\begin{bmatrix}\mathbf{x}\\ 1\end{bmatrix}$ $=H\begin{bmatrix}\mathbf{u}\\ 1\end{bmatrix}=\begin{bmatrix}RS&\boldsymbol{\mu}\\ \mathbf{0}&1\end{bmatrix}\begin{bmatrix}\mathbf{u}\\ 1\end{bmatrix},$
Figure 3. Integral of 3D Gaussian Density along a Ray. Under the transformation from the world frame to the canonical frame, the ray integral remains invariant. The green box highlights the product of density and ray segment length (see Eq. 5).
@@ 3.2. Exact Volumetric Gaussian Rendering
Figure 4. Pipeline Flowchart. Illustration of forward and backward gradient propagation from the Gaussian parameters $\mathbf{s}$, $\mathbf{q}$, $\boldsymbol{\mu}$, and the coefficient $\sigma$ to the transmittance.
To ensure that the transmittance contributed by each Gaussian along the ray is preserved across coordinate systems, we require that the elementary transmittance (i.e., the product of density and segment length) remains invariant under the transformation (see Figure 3):
(5) $\mathcal{G}_{\mathbf{I},\mathbf{0}}(\mathbf{u})\,\left|\Delta\mathbf{u}\right|=\mathcal{G}_{\Sigma,\boldsymbol{\mu}}(\mathbf{x})\,\left|\Delta\mathbf{x}\right|.$
(6) $\mathcal{G}_{\mathbf{I},\mathbf{0}}(\mathbf{u})=\frac{1}{\rho}\exp\left(-\frac{1}{2}\mathbf{u}^{\top}\mathbf{u}\right),$
This leads to an important conclusion that the exact closed-form solution for the transmittance integral of each Gaussian along the ray $\mathbf{r}(t)$ can thus be obtained by computing the accumulated density along the transformed ray $\mathbf{r}_{u}(t)=\mathbf{o}_{u}+t\mathbf{d}_{u}$ in the canonical isotropic Gaussian. This admits a closed-form solution in terms of the perpendicular distance from the ray to the Gaussian center, as captured in Theorem B.1. (see Appendix B for a complete proof):
[…]
(8) $\mathbf{d}_{u}=\left(S^{-1}R^{\top}\right)\mathbf{d}\ ,\quad\mathbf{m}_{u}=\mathbf{o}_{u}\times\mathbf{d}_{u},$
where $\mathbf{o}_{u}=\left(S^{-1}R^{\top}\right)\left(\mathbf{o}-\boldsymbol{\mu}\right)$ denotes the optical center in canonical space. Based on this, the square of $\mathrm{D}_{\mathbf{\mu},\Sigma}(\mathbf{o},\mathbf{d})$ can be efficiently calculated as:
(9) $\left(\mathrm{D}_{\mathbf{\mu},\Sigma}(\mathbf{o},\mathbf{d})\right)^{2}=\frac{\mathbf{m}_{u}^{\top}\mathbf{m}_{u}}{\mathbf{d}_{u}^{\top}\mathbf{d}_{u}}.$
By substituting this distance square into Eq. 7, and then the derived transmittance value into Eq. 3, we compute the accumulated color along the ray.
To minimize the computational overhead caused by automatic differentiation during training, we explicitly derive the ray-integral function and the gradients from the transmittance to all the Gaussian parameters, specifically $\frac{\partial T}{\partial\mathbf{s}},\frac{\partial T}{\partial\mathbf{q}},\frac{\partial T}{\partial\boldsymbol{\mu}}$. The detailed derivations for these derivatives are provided in Appendix C and visually explained in the Figure 4.
3.3. Efficient Ray-Particle Association via PBF
Given an exact volumetric Gaussian rendering method, both its final accuracy—particularly under large FoV camera models—and its runtime efficiency depend heavily on the ray-particle association strategy. To determine which Gaussians contribute to each ray, we simplify the association problem by mapping Camera CSFs to PBFs—the tight frustums enclosing each Gaussian at a certain distance from its center. A Gaussian is assigned to a CSF if its PBF intersects the CSF. All Gaussians associated with a CSF are then considered contributors to every ray within that CSF.

### v2

# 3DGEER: 3D Gaussian Rendering Made Exact and Efficient for Generic Cameras

@@ Abstract
3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) achieves an appealing balance between rendering quality and efficiency, but relies on approximating 3D Gaussians as 2D projections—an assumption that degrades accuracy, especially under generic large field-of-view (FoV) cameras. Despite recent extensions, no prior work has simultaneously achieved both projective exactness and real-time efficiency for general cameras. We introduce 3DGEER, a geometrically exact and efficient Gaussian rendering framework. From first principles, we derive a closed-form expression for integrating Gaussian density along a ray, enabling precise forward rendering and differentiable optimization under arbitrary camera models. To retain efficiency, we propose the Particle Bounding Frustum (PBF), which provides tight ray–Gaussian association without BVH traversal, and the Bipolar Equiangular Projection (BEAP), which unifies FoV representations, accelerates association, and improves reconstruction quality. Experiments on both pinhole and fisheye datasets show that 3DGEER outperforms prior methods across all metrics, runs $5\times$ faster than existing projective exact ray-based baselines, and generalizes to wider FoVs unseen during training—establishing a new state of the art in real-time radiance field rendering.
1 Introduction
@@
Initiated by 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023), volumetric particle rendering methods formulated under splatting (Yu et al., 2024a; Yan et al., 2024; Lu et al., 2024; Kheradmand et al., 2024; Charatan et al., 2024; Mallick et al., 2024) have gained significant popularity due to their impressive visual fidelity and high rendering speed. The core efficiency stems from EWA splatting (Zwicker et al., 2001), which uses a local affine approximation to project a 3D Gaussian as a 2D Gaussian on the image plane. While effective in narrow field-of-view (FoV) scenarios, this approximation introduces substantial errors in wide-FoV settings, where nonlinear distortions cause the true projection to deviate significantly from a symmetric 2D Gaussian—especially in fisheye images commonly encountered in robotics and autonomous driving (Rashed et al., 2021; Courbon et al., 2007; Kumar et al., 2023; Zhang et al., 2016; Yogamani et al., 2019; 2024; Sekkat et al., 2022). Although recent works have attempted to enable differentiable rendering under specific fisheye camera models (Liao et al., 2024) or more general camera models (Huang et al., 2024b), these approaches remain fundamentally constrained by their reliance on 2D projective-space approximations and have yet to yield an exact closed-form solution.
Meanwhile, Celarek et al. (2025) questions the necessity of exact rendering by scaling up the number of Gaussians (up to 1M) to mitigate projective approximation in splatting. However, the results are limited to small object-centric scenes. In large-scale scene reconstructions, the required number of Gaussians varies significantly across objects and frustums, making scaling up to larger scenes prohibitively costly. Further, we show projective errors under generic distorted cameras are substantially larger, where simply scaling up the Gaussian counts cannot close the quality gap (see Fig. 5).
@@
A parallel line of work tackles volumetric particle rendering directly in a ray-tracing/marching formulation (Yu et al., 2024b; Condor et al., 2025; Mai et al., 2025; Moënne-Loccoz et al., 2024; Talegaonkar et al., 2025; Gu et al., 2025). These approaches naturally support large-FoV camera models and are free from projective approximation error. However, achieving high efficiency remains a key challenge—particularly in associating each rendering ray with its most relevant 3D particles. To address this, significant effort has been devoted to accelerating it, such as adopting advanced Bounding Volume Hierarchies (BVH) employed in EVER (Mai et al., 2025) and 3DGRT (Moënne-Loccoz et al., 2024). Nevertheless, due to the inherent algorithmic complexity and difficulty in parallelization, pure ray-based methods have yet to match the efficiency of 3DGS. Alternatively, a hybrid approach has been proposed in 3DGUT (Wu et al., 2025) which adopts ray-tracing rendering with rasterization to solve ray-particle association efficiently. However, the inherited projective approximation error from the association stage still compromises rendering exactness in projective geometry and even risks grid-line artifacts (see Fig. 1 Left). To date, no existing method achieves both the geometric exactness of ray-based rendering and the high efficiency of splatting.
@@
In this work, we propose 3DGEER—a novel 3D Gaussian Exact and Efficient Rendering method that achieves exactness in projective geometry (see definition in Sec. A) while maintaining high efficiency comparable to splatting for generic cameras (both perspective and a variety of fisheye projection models). 3DGEER is built upon three key contributions: (i) We revisit the heuristic maximum response function (Keselman and Hebert, 2022) of Gaussians from first principles and reveal its connection with the projective exactness. Based on it, we further derive the differentiable rendering framework with numerical stability. (ii) To perform efficient ray-particle association without compromising geometric exactness, we reformulate the problem as associating Camera Sub-frustums (CSF) with Particle Bounding Frustums (PBF)—analogous to the tile-AABB mapping used in 3DGS. For each 3D Gaussian, we derive a closed-form solution to compute the PBF tightly bounding it, enabling highly efficient association. (iii) In addition, we propose to uniformly sample rays in a novel Bipolar Equiangular Projection (BEAP) space to apply color supervision. This ray sampling strategy not only aligns image-space partitioning and the underlying CSF but also improves conventional pinhole or equidistant projections in reconstruction quality.
Extensive experiments on both fisheye and pinhole datasets—including ScanNet++ (Yeshwanth et al., 2023), Zip-NeRF (Barron et al., 2023), Aria (Lv et al., 2025), and MipNeRF (Barron et al., 2021)—show that 3DGEER consistently outperforms prior and concurrent methods, achieving 0.9–4.8 PSNR gains under challenging wide-FoV conditions, and 0.3–2.0 on pinhole views. Remarkably, our approach even surpasses splat-based methods on wide-FoV cameras when trained only with narrow-FoV data, highlighting its strong generalization to extremely wide-FoV rendering.
Figure 1: Linear Approximation Error in Ray-Particle Association. (Left) Grid-line artifacts caused by inaccurate UT association. (Right) Diagram illustrating our association and comparison with others. Our method avoids the intermediate conic approximation by directly computing the exact bounding structure from the true 3D covariance.
2 Related Works
@@
Table 1: Summary of Particle Rendering Methods w.r.t. Projective Exactness. The top section lists splatting-based methods, while the bottom section presents ray-based approaches—some of which still rely on projective approximated association (e.g., EWA or Unscented Transform).
@@
Method Scope Type FoV FPS Meth.
3DGS (Kerbl et al., 2023) Splat. EWA Img. 3D Gauss.
FisheyeGS (Liao et al., 2024) Splat. EWA Img. 3D Gauss.
GS++  (Huang et al., 2024b) Splat. EWA Img. 3D Gauss.
2DGS (Huang et al., 2024a) Ray Splat Bound. Img. 2D Gauss.
GOF (Yu et al., 2024b) Ray EWA Img. 3D Gauss.
EVER (Mai et al., 2025) Ray BVH Scene 3D Ellip.
3DGRT (Moënne-Loccoz et al., 2024) Ray BVH Scene 3D Gauss.
3DGUT (Wu et al., 2025) Ray UT Img. 3D Gauss.
HTGS (Hahlbohm et al., 2025) Ray Splat Bound. Img. 3D Gauss.
@@
2.1 Splatting-Based Gaussian Rendering
Most 3DGS variants (Yu et al., 2024a; Zhang et al., 2024; Rota Bulò et al., 2024; Gao et al., 2025; Zhao et al., 2025; Kerbl et al., 2024; Hou et al., 2025; Liu et al., 2025) adopt the original EWA splatting strategy (Zwicker et al., 2001; Ren et al., 2002), projecting 3D Gaussians onto the image plane and approximating their appearance as 2D conics. To support differentiable rendering under large-FoV camera models with distortion, methods such as FisheyeGS (Liao et al., 2024) and GS++ (Huang et al., 2024b) estimate the Jacobian of equidistant or spherical projections to adjust the 2D mean and covariance of the projected Gaussians. However, these splatting-based approaches still rely on a first-order Taylor expansion of the highly non-linear projection function. In wide-FoV settings, higher-order terms are non-negligible, resulting in significant approximation errors and degraded reconstruction quality.
@@ 2.1 Splatting-Based Gaussian Rendering
2.2 Ray-Based Gaussian Rendering
Keselman and Hebert (2022) introduces an assumption that a ray intersects the Gaussian density function at the maximum intensity response. Building on this idea, 3DGRT (Moënne-Loccoz et al., 2024) adopts this heuristic intersection to establish a 3D Gaussian ray-tracing framework, which enables ray-space sampling without explicit projection and alleviates wide-FoV distortions. However, 3DGRT introduces substantial computational overhead, as the ray–particle association relies on costly scene-level spatial partitioning structures for BVHs, ultimately resulting in low frame rates.
Alternatively, EVER (Mai et al., 2025) replaces the Gaussian representation with simplified ellipsoids of constant density, enabling a closed-form solution for transmittance-weighted ray integration. However, this substitution reduces the expressive capacity of the representation compared to full 3D Gaussians and still suffers from limited runtime efficiency due to its reliance on BVH-based structures. 2DGS (Huang et al., 2024a) substitutes 3D Gaussians with 2D surfels, allowing exact ray–surface intersections. HTGS (Hahlbohm et al., 2025) reintroduces the full Gaussian expressions while adopting 2DGS’s ray–splat association. Yet, because its association remains restricted to screen space, it falls short of ensuring projective exactness under wider FoVs. Our contribution lies in deriving closed-form bounds and a full rendering pipeline directly in a camera-agnostic angular domain (see details in Sec. D) and we especially discuss the relation and differences in computation to prior 2DGS and HTGS parametrization and rendering in Sec. D.3.
@@ 2.2 Ray-Based Gaussian Rendering
3DGUT (Wu et al., 2025), a follow-up to 3DGRT, improves runtime efficiency by adopting splatting-based particle association, achieving frame rates comparable to 3DGS. However, its acceleration relies on the Unscented Transform (UT), a sampling-based approximation used to estimate 2D conics in projective image space. A similar limitation exists in GOF (Yu et al., 2024b), which directly applies the EWA trick to fit 2D bounding conics and projective centers for association. Notably, under extreme-FoV conditions, imprecise AABB estimation and the resulting inaccurate tile-to-AABB associations can lead to visible grid-line artifacts (see Fig. 1 Left).
@@
In contrast to these methods, our approach solves the ray-particle association problem directly at the frustum level in 3D space (see Fig. 1 Right), preserving the geometric exactness of ray-based rendering across arbitrary camera models while maintaining rendering speeds comparable to 3DGS. Tab. 1 summarizes representative methods that aim to address projective approximation errors, highlighting key differences across several important dimensions.
@@
Figure 2: Gradient Flow and Canonical Transformation. (Left) Illustration of forward and backward gradient propagation from the Gaussian parameters $\mathbf{s}$, $\mathbf{q}$, $\bm{\mu}$, and the opacity coefficient $\sigma$ to the transmittance. (Right) Integral of 3D Gaussian density along a ray, where the green box highlights the product of density and ray segment length (see Eq. B.5).
3 Method
We propose 3DGEER, a Gaussian rendering framework that achieves both geometric exactness and real-time efficiency under arbitrary FoV cameras. Our method addresses the two key bottlenecks of prior work—projection approximation in splatting-based rendering and BVH overhead in ray-based rendering—through three complementary components:
  • Ray-based rendering formulation (Sec. 3.1): a closed-form differentiable solution for integrating Gaussian density along 3D rays, eliminating linearization error from projection-based splatting.
  • Particle Bounding Frustum (PBF) (Sec. 3.2): a novel ray–particle association scheme that replaces BVH traversal with exact frustum–particle tests, yielding both accuracy and efficiency.
  • Bipolar Equiangular Projection (BEAP) (Sec. 3.3): an image representation unifying arbitrary FoV camera models, which accelerates PBF computation and improves reconstruction quality.
The Appendix further provides: (i) the necessary preliminaries and notations (Sec. B.1), (ii) the complete proof to Sec. 3.1 establishing the connection between maximum response and projective exactness (Sec. B), (iii) the full derivations of backward gradients (Sec. C) and its numerical stability, and (iv) the full derivations of geometrically exact association (Sec. D).
3.1 Rendering with Exact Projective Geometry
One key challenge in geometrically exact Gaussian rendering is to obtain a closed-form expression for integrating density and color along a ray. Our strategy is to map each anisotropic Gaussian in world space into a canonical coordinate system where all Gaussians become isotropic. This reduces the transmittance integral to a measure-preserving change of variables, admitting a simple closed-form solution.
Canonical transformation.
For each Gaussian with mean $\bm{\mu}$, rotation $R$, and scaling $S$, we define the mapping
$\begin{bmatrix}\mathbf{x}\\ 1\end{bmatrix}=\begin{bmatrix}RS&\bm{\mu}\\ \mathbf{0}&1\end{bmatrix}\begin{bmatrix}\mathbf{u}\\ 1\end{bmatrix}$, (1)
which maps a canonical coordinate $\mathbf{u}$ to world coordinates $\mathbf{x}$. This implies $\Delta\mathbf{x}=RS\Delta\mathbf{u}$ and absorbs the Jacobian determinant into the density measure (see Fig. 2 Right). Consequently, each Gaussian reduces to the shared isotropic form
$\mathcal{G}_{\mathbf{I},\mathbf{0}}(\mathbf{u})=\frac{1}{\rho}\exp\!\left(-\tfrac{1}{2}\|\mathbf{u}\|^{2}\right).$ (2)
Closed-form transmittance.
For a ray $\mathbf{r}(t)=\mathbf{o}+t\mathbf{d}$, its canonical form is $\mathbf{r}_{u}(t)=\mathbf{o}_{u}+t\mathbf{d}_{u}$ with
$\mathbf{o}_{u}=S^{-1}R^{\top}(\mathbf{o}-\bm{\mu}),\quad\mathbf{d}_{u}=S^{-1}R^{\top}\mathbf{d}.$ (3)
The accumulated transmittance is then
$T(\mathbf{o},\mathbf{d})=\sigma\int_{t\in\mathbb{R}}\mathcal{G}_{\mathbf{I},\mathbf{0}}(\mathbf{o}_{u}+t\mathbf{d}_{u})\,dt=\sigma\exp\!\left(-\tfrac{1}{2}D_{\mu,\Sigma}^{2}(\mathbf{o},\mathbf{d})\right),$ (4)
where
$D_{\mu,\Sigma}^{2}(\mathbf{o},\mathbf{d})=\frac{\|\mathbf{o}_{u}\times\mathbf{d}_{u}\|^{2}}{\|\mathbf{d}_{u}\|^{2}}$ (5)
is the squared perpendicular (Mahalanobis) distance from the ray to the Gaussian center.
Discussion.
This result reveals that the transmittance depends only on the ray’s distance to the Gaussian in the canonical space, yielding an efficient and geometrically exact closed-form solution. Interestingly, this expression is equivalent to the “maximum response” heuristic used in prior works (Keselman and Hebert, 2022; Moënne-Loccoz et al., 2024), while our mathematical derivation reveals its underlying projective exactness (see details in Sec. B). Furthermore, by following our closed-form derivation process and the corresponding backward gradient chain (see Fig. 2 Left and details in Sec. C), we explicitly avoid the numerical instabilities that prior ray-based work (Yu et al., 2024b) encounters and eliminates the need for redundant post-processing, such as filtering out degenerate Gaussians.